OpenStudy (anonymous):

A cylinder rod formed from silicon is 46.0 cm long and has a mass of 3.00 kg. The density of silicon is 2.33 g/cm^3. What is the diameter of the cylinder? (the volume of cylinder is given by (pi)r^2h, where r is the radius and h is the length)

OpenStudy (jfraser):

you know the mass of the cylinder and the density of the substance it's made of. Find the \(volume\) of the cylinder. \[Volume = \frac{mass}{density}\] You know the formula for the volume of a cylinder has radius in it. Find the \(radius\) of the cylinder\[Volume = \pi *r^2*h\]\[r = \sqrt{\frac{Volume}{\pi * h}}\]